# The T12 System for AudioMOS Challenge 2025: Audio Aesthetics Score Prediction System Using KAN- and VERSA-based Models

### v1

@@ Abstract
We propose an audio aesthetics score (AES) prediction system by CyberAgent (AESCA) for AudioMOS Challenge 2025 (AMC25) Track 2. The AESCA comprises a Kolmogorov–Arnold Network (KAN)-based audiobox aesthetics and a predictor from the metric scores using the VERSA toolkit. In the KAN-based predictor, we replaced each multi-layer perceptron layer in the baseline model with a group-rational KAN and trained the model with labeled and pseudo-labeled audio samples. The VERSA-based predictor was designed as a regression model using extreme gradient boosting, incorporating outputs from existing metrics. Both the KAN- and VERSA-based models predicted the AES, including the four evaluation axes. The final AES values were calculated using an ensemble model that combined four KAN-based models and a VERSA-based model. Our proposed T12 system yielded the best correlations among the submitted systems, in three axes at the utterance level, two axes at the system level, and the overall average.
@@ I Introduction
The field of generative artificial intelligence (AI) has witnessed significant advancements, particularly in the creation of synthetic audio encompassing speech, music, and other sounds. Because the primary users of these audio outputs are humans, evaluating the quality and effectiveness of these systems poses challenges similar to those found in the VoiceMOS challenge (VMC) for speech synthesis [huang2022vmc22, Cooper2023vmc23, Huang2024vmc24]. To engage with the latest developments in this field, the scope of the challenge was expanded and is now recognized as the AudioMOS Challenge 2025 (AMC25) [^1] https://sites.google.com/view/voicemos-challenge/audiomos-challenge-2025.
This challenge includes Track 2 that focused on assessing the outputs from text-to-speech (TTS), text-to-audio (TTA), and text-to-music (TTM) systems using the aesthetic scores (AES), a novel evaluation framework [Tjandra2024audiobox]. This framework aims to clarify evaluation by considering four distinct dimensions. Production Quality (PQ) examines technical attributes such as clarity and sound fidelity. Production Complexity (PC) assesses the intricacy of audio composition. Content Enjoyment (CE) evaluates subjective enjoyment and artistic value, whereas Content Usefulness (CU) considers the audio’s potential for creative projects.
@@ II Submitted T12 System (AESCA)
Figure 1 shows the proposed T12 system named “AESCA (AES prediction system by CyberAgent).” AESCA comprises a baseline system [Tjandra2024audiobox] extended by a Kolmogorov–Arnold network (KAN) [liu2025kan] and a VERSA-based predictor [shi2025versa]. In this section, we introduce the baseline model and provide details regarding each predictor and the ensemble model.
@@ II-A Baseline Model
Audiobox aesthetics [Tjandra2024audiobox] is a reference-free AES predictor designed to assess the AES of arbitrary audio clips, including speech and music at the utterance level. The model was built on a pretrained WavLM encoder comprising 12 transformer layers with a hidden size of 768 [Chen2022wavlm]. The encoder outputs branches into four independent paths, each of which applies learnable scalar weights to the outputs of each layer, aggregating information across the layers and time steps. These weighted audio representations were then fed into their respective multi-layer perceptrons (MLPs). They were responsible for regressing one of the following AES dimensions: PQ, PC, CE, or CU. The model was pretrained on a large dataset of approximately 500 hours, comprising 97,000 audio samples evenly distributed across speech, sound, and music.
@@ II-B KAN-based Audiobox Aesthetics
KAN augments neural networks by replacing fixed activation functions with learnable rational functions for each input–output pair, enabling highly expressive nonlinear modeling [liu2025kan]. For instance, the Kolmogorov–Arnold transformer (KAT) [yang2025kat] replaces MLP layers with group-rational KAN (GR-KAN) layers, and KAT improves the image recognition performances compared with the vision transformer [dosovitskiy2021vit]. Moreover, it is compatible with pretrained MLP weights, allowing seamless integration into existing transformer architectures without retraining, and offering improved flexibility and modeling capacity.
In the proposed KAN-based model, we replaced each MLP layer in the baseline model with a GR-KAN. This is expected to enable the model to effectively utilize the parameters of the baseline model [Tjandra2024audiobox], which was pretrained on a large amount of AES data while allowing for a more flexible representation.
@@ II-C VERSA-based predictor
AES correlates significantly with various evaluation metrics, indicating the overlapping yet complementary aspects of audio quality [Tjandra2024audiobox]. Thus, we considered a regression model that combines these metrics to support the KAN-based predictor.
The VERSA toolkit [shi2025versa] provides a range of reference-free evaluation models. For instance, VERSA includes DNSMOS [Reddy2021dnsmos], NISQA [Mittag2021nisqa], UTMOSv2 [Baba2024utmosv2], Torch-Squim, which predicts subjective quality scores (PESQ, STOI, and SI-SDR)[Kumar2023torchaudio], and the prompting audio-language model (PAM) for generative audio quality [Deshmukh2023pam]. VERSA also supports the audiobox aesthetics [Tjandra2024audiobox], which is the baseline model for this challenge.
Our VERSA-based predictor treats these evaluation metrics as input features and utilizes an extreme gradient boosting (XGBoost) [Chen2016xgboost] regression model to predict the AES.
@@ III-A Dataset
We used the AMC25 dataset for Track2222https://github.com/facebookresearch/audiobox-aesthetics/tree/main/audiomos2025_track2, including LibriTTS [zen19_interspeech], Common Voice 13.0 [ardila-etal-2020-common], EARS [richter2024ears], MUSDB18-HQ [musdb18-hq], AudioSet [Gemmeke2017audioset], and MusicCaps [agostinelli2023musiccaps]. The dataset contained 2,700 and 250 samples for the training and dev sets, respectively. Due to the absence of 70 samples in the dataset, the unavailable data were removed during the training phase.
In addition, to train the TTA and TTM data, we utilized the PAM dataset [Deshmukh2023pam], which contains both natural and generated sounds. Moreover, the data were labeled based on the AES listening tests [Tjandra2024audiobox] and are publicly available in the repository [^3] https://github.com/facebookresearch/audiobox-aesthetics/tree/main/evaluation_data. For training, we separated the PAM data into training and development sets, each with 900 and 100 samples, respectively, and added them to the AMC25 dataset. The main and out-of-domain tracks of the VMC 2022 dataset (VMC22) [huang2022vmc22], which includes natural, TTS, and voice conversion speech, were also utilized as unlabeled data for training the TTS data. From the VMC22 dataset, 8,459 samples were used, with pseudo-labels generated based on the strategy described in the next section.
@@ III-B Experimental Conditions
For the KAN-based model, we set the batch size to 40, number of epochs to 10, and learning rate to $1\times 10^{-4}$ with a schedule-free version of AdamW as the optimizer [Defazio2024schedulefree]. The model was trained for approximately 6–8 GPU hours on an NVIDIA A100 GPU with 80 GB memory.
For the VERSA-based predictor, we used 28 outputs from existing metrics [^4] As VERSA is an actively maintained toolkit, we used available scalar-valued and reference-free metrics in the codebase at commit c2a7663. as input for the regression model. The outputs of the regressor correspond to PQ, PC, CE, and CU. We trained each model using 10-fold cross-validation. We used Optuna [optuna] to search for the hyperparameters to minimize the MSE of each target.
@@ III-C Training Strategies
For the KAN-based model, we employed a joint strategy of supervised and semi-supervised training using an iterative pseudo-labeling (IPL) approach inspired by [Xu2020ipl, Hwang2022nst]. Initially, pseudo-labels were generated using the baseline model as the “teacher.” These pseudo-labels were combined with the AMC25 and PAM datasets, which contained the labeled data. The “student” model is then fine-tuned using this dataset. The student replaces the teacher model if the developmental loss is low. The process halts when no improvement occurs in the development loss. In our experiment, we set the maximum number of updates to five.

### v2

@@ Abstract
We propose an audio aesthetics score (AES) prediction system by CyberAgent (AESCA) for AudioMOS Challenge 2025 (AMC25) Track 2. The AESCA comprises a Kolmogorov–Arnold Network (KAN)-based audiobox aesthetics and a predictor from the metric scores using the VERSA toolkit. In the KAN-based predictor, we replaced each multi-layer perceptron layer in the baseline model with a group-rational KAN and trained the model with labeled and pseudo-labeled audio samples. The VERSA-based predictor was designed as a regression model using extreme gradient boosting, incorporating outputs from existing metrics. Both the KAN- and VERSA-based models predicted the AES, including the four evaluation axes. The final AES values were calculated using an ensemble model that combined four KAN-based models and a VERSA-based model. Our proposed T12 system yielded the best correlations among the submitted systems, in three axes at the utterance level, two axes at the system level, and the overall average. We also released the inference model of the proposed KAN-based predictor (KAN #1-#4)  [^1] https://github.com/CyberAgentAILab/aesca.
@@ I Introduction
The field of generative artificial intelligence (AI) has witnessed significant advancements, particularly in the creation of synthetic audio encompassing speech, music, and other sounds. Because the primary users of these audio outputs are humans, evaluating the quality and effectiveness of these systems poses challenges similar to those found in the VoiceMOS challenge (VMC) for speech synthesis [12, 7, 13]. To engage with the latest developments in this field, the scope of the challenge was expanded and is now recognized as the AudioMOS Challenge 2025 (AMC25) [^2] https://sites.google.com/view/voicemos-challenge/audiomos-challenge-2025.
This challenge includes Track 2 that focused on assessing the outputs from text-to-speech (TTS), text-to-audio (TTA), and text-to-music (TTM) systems using the aesthetic scores (AES), a novel evaluation framework [22]. This framework aims to clarify evaluation by considering four distinct dimensions. Production Quality (PQ) examines technical attributes such as clarity and sound fidelity. Production Complexity (PC) assesses the intricacy of audio composition. Content Enjoyment (CE) evaluates subjective enjoyment and artistic value, whereas Content Usefulness (CU) considers the audio’s potential for creative projects.
@@ II Submitted T12 System (AESCA)
Figure 1 shows the proposed T12 system named “AESCA (AES prediction system by CyberAgent).” AESCA comprises a baseline system [22] extended by a Kolmogorov–Arnold network (KAN) [16] and a VERSA-based predictor [21]. In this section, we introduce the baseline model and provide details regarding each predictor and the ensemble model.
@@ II-A Baseline Model
Audiobox aesthetics [22] is a reference-free AES predictor designed to assess the AES of arbitrary audio clips, including speech and music at the utterance level. The model was built on a pretrained WavLM encoder comprising 12 transformer layers with a hidden size of 768 [5]. The encoder outputs branches into four independent paths, each of which applies learnable scalar weights to the outputs of each layer, aggregating information across the layers and time steps. These weighted audio representations were then fed into their respective multi-layer perceptrons (MLPs). They were responsible for regressing one of the following AES dimensions: PQ, PC, CE, or CU. The model was pretrained on a large dataset of approximately 500 hours, comprising 97,000 audio samples evenly distributed across speech, sound, and music.
@@ II-B KAN-based Audiobox Aesthetics
KAN augments neural networks by replacing fixed activation functions with learnable rational functions for each input–output pair, enabling highly expressive nonlinear modeling [16]. For instance, the Kolmogorov–Arnold transformer (KAT) [24] replaces MLP layers with group-rational KAN (GR-KAN) layers, and KAT improves the image recognition performances compared with the vision transformer [10]. Moreover, it is compatible with pretrained MLP weights, allowing seamless integration into existing transformer architectures without retraining, and offering improved flexibility and modeling capacity.
In the proposed KAN-based model, we replaced each MLP layer in the baseline model with a GR-KAN. This is expected to enable the model to effectively utilize the parameters of the baseline model [22], which was pretrained on a large amount of AES data while allowing for a more flexible representation.
@@ II-C VERSA-based predictor
AES correlates significantly with various evaluation metrics, indicating the overlapping yet complementary aspects of audio quality [22]. Thus, we considered a regression model that combines these metrics to support the KAN-based predictor.
The VERSA toolkit [21] provides a range of reference-free evaluation models. For instance, VERSA includes DNSMOS [19], NISQA [17], UTMOSv2 [4], Torch-Squim, which predicts subjective quality scores (PESQ, STOI, and SI-SDR)[15], and the prompting audio-language model (PAM) for generative audio quality [9]. VERSA also supports the audiobox aesthetics [22], which is the baseline model for this challenge.
Our VERSA-based predictor treats these evaluation metrics as input features and utilizes an extreme gradient boosting (XGBoost) [6] regression model to predict the AES.
@@ III-A Dataset
We used the AMC25 dataset for Track2333https://github.com/facebookresearch/audiobox-aesthetics/tree/main/audiomos2025_track2, including LibriTTS [25], Common Voice 13.0 [3], EARS [20], MUSDB18-HQ [18], AudioSet [11], and MusicCaps [1]. The dataset contained 2,700 and 250 samples for the training and dev sets, respectively. Due to the absence of 70 samples in the dataset, the unavailable data were removed during the training phase.
In addition, to train the TTA and TTM data, we utilized the PAM dataset [9], which contains both natural and generated sounds. Moreover, the data were labeled based on the AES listening tests [22] and are publicly available in the repository [^4] https://github.com/facebookresearch/audiobox-aesthetics/tree/main/evaluation_data. For training, we separated the PAM data into training and development sets, each with 900 and 100 samples, respectively, and added them to the AMC25 dataset. The main and out-of-domain tracks of the VMC 2022 dataset (VMC22) [12], which includes natural, TTS, and voice conversion speech, were also utilized as unlabeled data for training the TTS data. From the VMC22 dataset, 8,459 samples were used, with pseudo-labels generated based on the strategy described in the next section.
@@ III-B Experimental Conditions
For the KAN-based model, we set the batch size to 40, number of epochs to 10, and learning rate to $1\times 10^{-4}$ with a schedule-free version of AdamW as the optimizer [8]. The model was trained for approximately 6–8 GPU hours on an NVIDIA A100 GPU with 80 GB memory.
For the VERSA-based predictor, we used 28 outputs from existing metrics [^5] As VERSA is an actively maintained toolkit, we used available scalar-valued and reference-free metrics in the codebase at commit c2a7663. as input for the regression model. The outputs of the regressor correspond to PQ, PC, CE, and CU. We trained each model using 10-fold cross-validation. We used Optuna [2] to search for the hyperparameters to minimize the MSE of each target.
@@ III-C Training Strategies
For the KAN-based model, we employed a joint strategy of supervised and semi-supervised training using an iterative pseudo-labeling (IPL) approach inspired by [23, 14]. Initially, pseudo-labels were generated using the baseline model as the “teacher.” These pseudo-labels were combined with the AMC25 and PAM datasets, which contained the labeled data. The “student” model is then fine-tuned using this dataset. The student replaces the teacher model if the developmental loss is low. The process halts when no improvement occurs in the development loss. In our experiment, we set the maximum number of updates to five.